### v4

# SATURN: SAT-based Reinforcement Learning to Unleash Language Model Reasoning

Abstract
How to design reinforcement learning (RL) tasks that effectively unleash the reasoning capability of large language models (LLMs) remains an open question. Existing RL tasks (e.g., math, programming, and constructing reasoning tasks) face three key limitations: ❶ Scalability. They rely heavily on human annotation or expensive LLM synthesis to generate sufficient training data. ❷ Verifiability. LLMs’ outputs are hard to verify automatically and reliably. ❸ Controllable Difficulty. Most tasks lack fine-grained difficulty control, making it challenging to train LLMs from easy to hard and progressively develop reasoning capability.
To address these limitations, we propose SATURN, a SAT-based RL framework that uses Boolean Satisfiability (SAT) problems to train and evaluate LLM reasoning. SATURN enables scalable task construction, rule-based verification, and precise difficulty control. SATURN designs a curriculum learning pipeline that continuously improves LLMs’ reasoning capability by constructing SAT tasks of increasing difficulty and training LLMs from easy to hard. To ensure stable training, we design a principled mechanism to control difficulty transitions.
We introduce SATURN-2.6k, a dataset of 2,660 SAT problems with varying difficulty. It supports the evaluation of how LLM reasoning changes with problem difficulty. We apply SATURN to DeepSeek-R1-Distill-Qwen and obtain SATURN-1.5B and SATURN-7B. We achieve several notable results: ❶ On SAT problems, SATURN-1.5B and SATURN-7B achieve average pass@3 improvements of +14.0 and +28.1, respectively. ❷ On math and programming tasks, SATURN-1.5B and SATURN-7B improve average scores by +4.9 and +1.8 on benchmarks (e.g., AIME, LiveCodeBench). ❸ Compared to the state-of-the-art (SOTA) approach in constructing RL tasks, SATURN achieves further improvements of +8.8%. We release the source code, data, and models to support future research at https://github.com/gtxygyzb/Saturn-code.
1 Introduction
Recently, reinforcement learning (RL) has become a promising paradigm for unleashing the reasoning capability of large language models (LLMs), particularly in math, programming, and logical reasoning (e.g., OpenAI-o1 [21], DeepSeek-R1 [11], Kimi-k1.5 [37]). During the RL training process, the design of RL tasks plays a critical role [14, 32, 31]. A well-designed RL task should elicit LLMs’ reasoning capability, fostering behaviors such as hesitation, reflection, backtracking, summarization, and verification [43, 44, 53, 31, 28].
However, how to design RL tasks that can continuously enhance LLMs’ reasoning capability remains an open question. We think a well-designed RL task for reasoning should satisfy the following three criteria: ❶ Scalability. RL training requires large-scale data. RL tasks should support scalable data without human annotation or expensive LLMs’ synthesis. ❷ Verifiability. RL rewards must be unambiguously correct. The outputs of LLMs for the task should be easy to verify. ❸ Controllable Difficulty. Reasoning capability emerges progressively [42]. RL tasks should support the difficulty control to enable curriculum learning, allowing LLMs to gradually develop more complex reasoning skills [16].
Table 1: The comparison between existing RL tasks and SATURN.
Tasks Scalability Verifiability Controllable Difficulty
ScaleQuest [12]
GSM8K (Math) [9]
LiveCodeBench [22]
Game Werewolf [45, 48]
LMRL Gym [2]
SPAG [7]
Knights&Knaves [46]
SATURN (Ours)
Table 1 shows the features of current mainstream RL tasks. None of them satisfy all three criteria. Existing RL tasks can be divided into two categories: (1) One category of RL tasks requires LLMs to solve math or programming problems, with rewards based on the correctness of the final answer or code [6, 27, 5, 9]. However, these tasks rely on human annotation for ground-truth solutions or test cases, suffer from a lack of high-quality problems, and offer only coarse control over reasoning difficulty [23, 26]. (2) Another category focuses on manually designed reasoning tasks [12, 45, 48, 46]. For instance, Logic-RL [47] leverages natural language logic K&K puzzles to improve LLMs’ reasoning capability through RL. However, they also present limitations, such as hard to scale up due to reliance on sampling from LLMs [7, 12], hard to verify despite relying on LLMs for cross-validation [51, 12, 41], and hard to control difficulty [45, 48].
In this paper, we aim to answer the following research question:
Key Question Can we design an RL task that supports scalability, verifiability, controllable difficulty, and enhances the reasoning capability of LLMs?
To this end, we propose Boolean Satisfiability (SAT) problem as the task for RL. Figure 1 shows an illustration of SAT problems and corresponding features. SAT satisfies all three desiderata we outlined earlier: ❶Scalability. SAT instances can be generated programmatically at scale without human annotation or LLM synthesis, allowing for virtually unlimited training data. ❷ Verifiability. SAT is a well-established NP-complete problem in theoretical computer science [10]. The correctness of a solution can be easily verified in linear time. But solving SAT problems requires complex reasoning. ❸ Controllable Difficulty. The difficulty of SAT instances can be precisely adjusted (e.g., number of variables, clauses), making it suitable for curriculum learning. What’s more, SAT serves as a universal substrate for limited forms of logical reasoning, as many problems in propositional logic, finite-domain first-order logic, and modal logic can be systematically reduced to SAT [17, 36, 24].
Figure 1: An illustration of SAT problems and its corresponding features.
Building on these advantages, we propose SAT-based reinforcement learning to Unleash LLMs ReasoNing, or SATURN. SATURN is a multi-stage curriculum learning-based RL framework that continuously improves the reasoning capability of LLMs. SATURN efficiently constructs SAT tasks with controllable difficulty and organizes them into progressive stages from easy to hard, allowing LLMs to develop reasoning skills step by step. To ensure training stability and effective progression, we design a principled mechanism to control difficulty transitions based on LLMs’ performance. SATURN enables smooth curriculum advancement and steady enhancement of reasoning capability.
We introduce the SATURN-2.6k dataset, consisting of 1,500 training instances, 160 test instances at the same difficulty as the training set, and 1,000 test instances from 10 harder unseen difficulty levels. The test set serves as a benchmark for systematically evaluating how LLMs’ reasoning capability varies with increasing SAT task difficulty. We release SAT construction scripts alongside the dataset, which enable the creation of virtually unlimited SAT instances.
We apply SATURN to DeepSeek-R1-Distill-Qwen-1.5B and 7B [11], obtaining SATURN-1.5B and SATURN-7B. Experiments show that SATURN effectively enhances LLMs’ reasoning capability in generalizable scenarios:
  • SATURN-1.5B and SATURN-7B achieve strong performance on SATURN-2.6k benchmarks. On unseen harder test set, two models achieve pass@3 improvements of +14.0 and +28.1 respectively.
  • The reasoning capability learned from SATURN transfers well to math and programming tasks, bringing average improvements of +4.9 and +1.8 on popular benchmarks such as AIME [3], AMC [4], MATH-500 [19], GPQA Diamond [35], and LiveCodebench [22] for two SATURN models.
  • Compared to the prior SOTA approach (e.g., Logic-RL), SATURN achieve average improvements of +8.8% on math and programming tasks.
2 SATURN
Figure 2: The overall framework of SATURN. It alternates between two interconnected loops: (1) Curriculum Estimation Loop. (2) LLMs Training Loop. The two loops iterate until the maximum number of curriculum stages is reached.
2.1 SATURN Learning Loop Framework
We introduce SATURN, a multi-stage RL framework that leverages SAT tasks to unleash LLMs’ reasoning via curriculum learning. As illustrated in Figure 2, SATURN alternates between two interconnected loops: Curriculum Estimation Loop dynamically constructs SAT instances of adjustable difficulty and evaluates LLMs’ performance to determine whether to advance the curriculum stage; LLMs Training Loop employs RL to optimize LLMs on current difficulty SAT tasks. The curriculum loop presented in Algorithm 1 proceeds as follows:
Step 1: Curriculum Estimation Loop. Given initial SAT difficulty, SAT_Construction generates a validation set of SAT instances. The LLM is evaluated on this set using the pass@1 metric. If the performance exceeds a predefined threshold $\epsilon$, the curriculum controller advances to a harder configuration with an increased estimated difficulty. Otherwise, SATURN process enters Step 2 LLMs training loop at the current SAT difficulty. This adaptive loop ensures that the LLM is always trained at the frontier of its reasoning capability, neither too easy nor too hard.
Step 2: LLMs Training Loop. For the current difficulty, SAT_Construction generates a set of training instances that are different from the validation set. These samples are then used to train LLMs with GRPO. The reward function encourages outputs that are both logically correct and properly formatted. The training loop proceeds until $\texttt{pass@1}>\epsilon$ on the validation set. After that, the process backs to Step 1 to reassess and potentially advance the curriculum.
The two loops iterate jointly. SATURN process terminates when a predefined total number of iterations is reached. Importantly, SATURN is not designed to replace math or programming tasks, but to serve as a complementary strategy for enhancing LLMs’ reasoning. In practice, SATURN can be integrated with math and programming tasks to enable a stronger training framework.
SATURN learning loop raises three core challenges: ❶ Section 2.2 introduces how to construct scalable and controllable SAT instances. ❷ Section 2.3 presents how to estimate instance difficulty for curriculum learning. ❸ Section 2.4 explains how to train LLMs on SAT tasks with RL.
2.2 SAT Instances Construction
In this subsection, we formalize the construction of SAT instances. A SAT problem determines whether a propositional formula can be satisfied by a Boolean truth assignment. Formally, we define a $(n,k,l)$-SAT instance in conjunctive normal form (CNF) as:
$\left\{\begin{aligned} \Phi=\ &\left(x_{a_{1,1}}\lor\neg x_{a_{1,2}}\lor\dots\lor x_{a_{1,n}}\right)\land\cdots\land\left(x_{a_{l,1}}\lor\dots\lor\neg x_{a_{l,n}}\right)\\ \text{where }&a_{i,j}\in\{1,\dots,k\},\quad i\in[1,l]_{\mathbb{Z}},j\in[1,n]_{\mathbb{Z}}\end{aligned}\right$. (1)
where each clause contains exactly $n$ variables (literals), each being either $x_{i}$ or its negation $\neg x_{i}$, $k$ is the total number of variables, and $l$ is the total number of clauses. Based on the definition, we design a SAT instance constructor, SAT_Construction(n,k,l,m)(n,k,l,m), which uniformly samples $m$ SAT instances from the space of $(n,k,l)$-SAT. By adjusting the parameters $(n,k,l,m)$, SAT_Construction enables the scalable and controllable construction of SAT instances. The design details of the constructor algorithm are provided in Appendix B. All generated SAT instances are guaranteed to be satisfiable.
2.3 Estimation of Task Difficulty
In this subsection, we present the estimation of SAT task difficulty for LLMs. This estimation also serves as the foundation for curriculum learning in LLMs.
As a canonical NP-complete problem [10], SAT admits a polynomial-time reduction from any other NP problem [25]. SAT exhibits a known phase transition phenomenon: when the clause-to-variable ratio $\alpha_{c}=l/k$ approaches a critical threshold (typically near $4.26$ for 3-SAT), the probability of satisfiability drops sharply, and problem difficulty peaks. This phenomenon probably stems from replica symmetry breaking (RSB) [54]: near $\alpha_{c}$, the solution space fractures into disconnected clusters separated by energy barriers. Beyond $\alpha_{c}$, the space collapses, reducing complexity.
However, RSB theory is designed for heuristic SAT solvers. For humans or LLMs solving SAT problems through logical steps such as trial, verification, and reasoning, such solver-like phase transitions are hardly observable in human-like thinking processes. While any $n$-SAT ($n>3$) can be reduced to 3-SAT [25], they differ significantly for LLMs in terms of solution space size and token length.
Prior work [18] on SAT tasks for LLMs typically categorized difficulty based on phase transition points. To systematically estimate task difficulty, we define an analytical estimator of the expected solution space size. Given a $(n,k,l)$ SAT instance, its difficulty for LLMs can be approximately estimated by:
$\footnotesize D(n,k,l)=\log_{2}(k)+2\log_{2}(l)-n+\frac{k}{n}$ (2)
Eq. (2) provides a more controllable, fine-grained estimation of SAT task difficulty. The detailed derivation is provided in Appendix C. To further validate Eq. (2), we evaluate LLMs’ performance on SAT instances with varying difficulty levels. As shown in Figure 3, each point represents a LLM’s average pass@3 on the same estimated difficulty instances. Pass@3 generally decreases as $D(n,k,l)$ increases, suggesting that our estimation aligns with the solvability trends observed in practical LLMs. Stronger LLMs maintain higher pass@3, while weaker LLMs exhibit lower scores overall. The validity of the estimation in Eq. (2) is further confirmed by ablation experiments, as detailed in Appendix C.
Figure 3: Scatter plots of pass@3 versus estimated difficulty $D(n,k,l)$ for different LLMs, with linear regression fits. The linear regression for two models achieve $R^{2}$ values of 0.707 and 0.724 respectively, suggesting a reasonably strong linear relationship between difficulty and pass@3.
2.4 Reinforcement Learning with GRPO
In this subsection, we introduce the single-stage RL training for given $(n,k,l)$-difficulty tasks. RL can further improve LLMs’ generalization by directly optimizing policy gradients over diverse reasoning trajectories [11]. Given the SAT tasks, we then train LLMs using the original sample-level GRPO to optimize the policy $\pi_{\theta}$ with KL divergence penalty. The GRPO objective function is defined as:
$\mathcal{L}_{\text{GRPO}}(\theta)=$ $\mathbb{E}{[q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(O|q)]}$ (3)
$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[r_{i,t}(\theta)\hat{A}_{i,t},\text{clip}\left(r_{i,t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right]-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\parallel\pi_{\text{ref}}\right]\right\}$
$\text{where }r_{i,t}(\theta)$ $=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})},\qquad\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(\mathbf{r})}{\text{std}(\mathbf{r})}$
[…]
3 Experiments
We apply SATURN to DeepSeek-R1-Distill-Qwen-1.5B and 7B, obtaining SATURN-1.5B and SATURN-7B. To evaluate the effectiveness of SATURN, we conduct a large-scale study to evaluate both models. In this section, we introduce our research questions (RQs), benchmarks, baselines, and evaluation metrics. For each RQ, we present the corresponding experimental design, results, and analysis in separate subsections.
3.1 Research Questions
Our study aims to answer the following research questions:
RQ1: How much improvement does SATURN achieve in solving SAT tasks? We evaluate SATURN-1.5B and SATURN-7B performance on SATURN-2.6k with different difficulty levels.
RQ2: How effectively does SATURN generalize to math and programming tasks? To evaluate the transferability of reasoning capabilities learned by SATURN, we evaluate the performance of LLMs on math and programming benchmarks and compare them with current SOTA LLMs.
RQ3: How does SATURN compare to prior RL tasks? To explore the relationship between SATURN and existing RL tasks, RQ3 investigates whether SATURN can (1) serve as a complementary task to math and programming, and (2) outperform other constructing RL tasks.
RQ4: How does SATURN affect LLMs reasoning trajectory? RQ4 explores whether SATURN influences the reasoning patterns of LLMs, particularly in terms of response length and the capability of verification. We investigate whether the reasoning improvements observed in SAT tasks generalize to math and programming.
3.2 Experimental Setup
SATURN Hyperparameters
For SATURN-1.5B and SATURN-7B, we set the initial SAT instance parameters $(n,k,l)$ to $(3,5,5)$ and $(3,5,13)$, respectively. In Curriculum Estimation Loop, the $\epsilon$ threshold is set to 0.5 for the 1.5B model and 0.75 for the 7B model. In LLMs Training Loop, we evaluate the pass@k with a step size of 250 training samples. The total number of curriculum iterations is set to 2. Detailed hyperparameters are provided in Appendix A. Ablation studies in Appendix H demonstrate the necessity of curriculum learning and the effectiveness of hyperparameters on SAT difficulty, thresholds, step sizes, etc.
Benchmarks.
❶Building upon SAT_Construction tool and difficulty estimation, we release SATURN-2.6k, a curated benchmark designed to evaluate LLMs’ reasoning capability across varying complexity. SATURN-2.6k consists of 1,500 training instances and 160 test instances that share the same estimated difficulty level. To assess performance under increasing task complexity, SATURN-2.6k further includes 1,000 test instances from 10 unseen harder difficulty levels, with 100 instances per level. These levels are selected based on our difficulty estimation $D(n,k,l)$, enabling a systematic analysis of how LLM performance changes as problem difficulty increases. Additionally, custom datasets of desired difficulty can be constructed using our open-sourced SAT_Construction tool. ❷ For math and programming tasks, following DeepSeek-AI [11], we use AIME 24/25 [3], AMC 22/23 [4], MATH-500 [19], GPQA Diamond [35], and LiveCodeBench v4_v5 subset [22].
Baseline Model.
We conduct evaluations against several 1.5B and 7B parameter reasoning models as the baselines, which include DeepSeek-R1-Distill-Qwen-1.5B & 7B [11], Still-3-1.5B-Preview [40], s1.1-1.5B & 7B [31], z1-7B [50], OpenThinker-7B [38], and DeepScaleR-1.5B-Preview [29]. In addition, we include a supervised fine-tuning (SFT)-only baseline trained on the Math training dataset [19], which provides step-by-step problem reasoning trajectories. We randomly select the most difficult Level-5 1,000 problems from training set for one epoch of SFT, following the same training template as DeepSeek-R1-Distill-Qwen. With the same dataset size, our setup enables a fair comparison between SFT and RL on SAT tasks.
Evaluation Metrics.
Following DeepSeek-AI [11], we use pass@k as the evaluation metric. Pass@k assesses the probability that at least one correct solution is generated within $k$ attempts. For SAT problems, we evaluate $\texttt{pass@k}\in\{1,3,5,7,10\}$ and sample 12 times per problem. For math and programming benchmarks, we use pass@1, following a context length of 32,768 and temperature = 0.6. More evaluation hyperparameters are provided in Appendix E. All experiments are conducted on NVIDIA 8×A100 (40GB) GPUs. Specific prompts are detailed in Appendix F.
3.3 RQ1: SATURN Substantially Improves Performance on SAT Tasks
We evaluate the performance of SATURN-1.5B and SATURN-7B on SAT tasks using SATURN-2.6k test set. Specifically, the evaluation involves unseen SAT instances that were not included in the training data. The results, presented in Table 2, 1014, and detailed in Appendix G, demonstrate the performance of LLMs across different SAT difficulties.
Table 2: Performance (pass@k, in %) on SATURN-2.6k test set across different difficulty levels.
Model SAT-(3,5,5) SAT-(3,5,8)
@1 @3 @5 @7 @10 @1 @3 @5 @7 @10
DeepSeek-R1-Distill-Qwen-1.5B 36.7 71.7 85.4 91.7 96.2 20.3 47.6 63.6 73.4 81.9
SATURN-1.5B-Iteration-1 59.7 90.4 97.1 99.1 99.8 41.0 74.0 85.6 91.1 95.6
SATURN-1.5B-Iteration-2 70.3 95.9 99.0 99.7 99.9 47.0 82.6 93.9 98.0 99.8
Model SAT-(3,5,13) SAT-(3,5,15)
@1 @3 @5 @7 @10 @1 @3 @5 @7 @10
DeepSeek-R1-Distill-Qwen-7B 53.9 86.2 94.2 97.3 99.3 39.3 74.9 88.3 94.3 98.3
SATURN-7B-Iteration-1 73.0 96.1 98.9 99.7 99.9 65.7 91.8 96.8 98.7 99.7
SATURN-7B-Iteration-2 89.5 99.0 99.9 100.0 100.0 85.4 98.3 99.8 99.9 100.0
SATURN substantially improves LLM performance on SAT tasks across varying difficulty levels. On the difficulty SAT-(3,5,5), SATURN-1.5B improves pass@1 from 36.7 to 59.7 at Iteration-1, and further to 70.3 at Iteration-2, achieving a total gain of +33.6. On the unseen harder test set (Table 11), SATURN-1.5B improves average pass@3 from 10.1 to 24.2, while SATURN-7B improves from 36.1 to 64.2. On average, these models achieve pass@3 improvements of +14.0 and +28.1 respectively, confirming that SATURN effectively enhances LLM reasoning across both seen and unseen SAT difficulties.
3.4 RQ2: SATURN Demonstrates Strong Generalization to Math and Programming
We assess whether the reasoning capability learned by SATURN generalizes to math and programming tasks. We evaluate SATURN-1.5B and SATURN-7B on a range of reasoning benchmarks. The results shown in Table 3 provide a detailed comparison.
Table 3: Performance comparison on math and programming Benchmarks
Model AIME 24/25 AMC 22/23 Math500 GPQA-D LiveCodeBench Avg.
GPT-4o (Aug’24) 11.7 - 79.5 52.1 31.7 -
Claude 3.5 Sonnet (Oct ’24) 15.7 - 77.1 59.9 38.1 -
s1.1-1.5B 1.7 25.3 42.2 29.3 2.2 20.1
Still-3-1.5B-Preview 23.3 74.7 84.6 34.8 17.1 46.9
DeepSeek-R1-Distill-Qwen-1.5B 21.6 65.1 83.6 30.3 16.4 43.4
   + SFT 25.0 68.7 82.0 34.3 14.6 44.9
SATURN-1.5B 28.3 73.5 84.6 37.4 17.4 48.2
z1-7B 8.3 39.8 74.2 35.4 19.3 35.4
s1.1-7B 21.7 61.4 80.8 43.4 12.8 44.0
OpenThinker-7B 26.7 53.0 88.6 42.9 21.5 46.5
DeepSeek-R1-Distill-Qwen-7B 50.0 80.7 93.2 49.0 35.4 61.7
SATURN-7B 48.3 85.5 95.0 50.5 37.7 63.4
❶SATURN shows strong generalization to math and programming tasks. On the AIME 24/25 benchmark, SATURN-1.5B outperforms z1-7B by 8.3 and s1.1-7B by 21.7. Similarly, SATURN-7B achieves a strong improvement on the Math500 dataset, increasing from 93.2 to 95.0. On LiveCodeBench, it improves from 35.4 to 37.7. On average, SATURN-1.5B improves by +4.9, and SATURN-7B improves by +1.8 across these benchmarks. These results highlight that SATURN enhances the reasoning performance of LLMs across various math and programming tasks, demonstrating strong generalization of the learned reasoning capabilities from SAT.
❷SATURN outperforms SFT on broader benchmarks. Consistent with the observations in SFT Memorizes, RL Generalizes [8], SFT improves performance on math-focused benchmarks (AIME, AMC, and Math500) that are similar to its supervised training domain. However, on LiveCodeBench, SFT drops from 16.4 to 14.6, exhibiting an alignment tax [33], where specializing on a narrow domain compromises performance on other tasks. In contrast, SATURN improves performance across all benchmarks, with SATURN-1.5B reaching 17.4 on LiveCodeBench. Averaging across all benchmarks, SATURN-1.5B outperforms the SFT counterpart by 3.3, demonstrating that SATURN generalizes effectively.
3.5 RQ3: SATURN Serves as a Complement and Further Enhances LLM Reasoning
RQ3 studies the relationship between SATURN and existing RL tasks. Beyond the DeepSeek-R1-Distill-Qwen-7B, we introduce two additional models: Qwen2.5-7B-Instruct-1M [49, 39] following Logic-RL [47] settings, and DeepScaleR-1.5B-Preview [29], which is further RL trained on 40k math and programming examples from DeepSeek-R1-Distill-Qwen-7B. We compare SATURN against several prior constructing RL task approaches, including Logic-RL [47], SPGA [7], and ScaleQuest [12], which represent strong baselines. Each approach is applied to different models for comparison. Results are summarized in Table 4.
Table 4: Comparison of SATURN and prior approaches across various LLMs.
Model AIME 24/25 AMC 22/23 Math500 GPQA-D LiveCodeBench Avg.
SPGA-3 (82k) 0.0 3.6 7.2 24.7 0.0 7.1
ScaleQuest (25k) 6.7 45.8 74.6 31.3 7.9 33.3
Qwen2.5-7B-Instruct-1M 5.0 41.0 74.4 32.3 9.8 32.5
   + Logic-RL (5k) 6.7 49.4 72.0 29.3 9.0 33.3
    + Saturn (1k) 10.0 47.0 74.8 37.9 11.3 36.2
DeepSeek-R1-Distill-Qwen-7B 50.0 80.7 93.2 49.0 35.4 61.7
   + Logic-RL (5k) 50.0 80.7 93.4 52.0 35.7 62.4
    + Saturn (1k) 48.3 85.5 95.0 50.5 37.7 63.4
DeepScaleR-1.5B-Preview 30.0 74.7 87.8 37.4 19.8 49.9
   + Logic-RL (5k) 28.3 77.1 86.4 35.9 20.7 49.7
    + Saturn (0.5k) 35.0 73.5 88.6 43.4 21.0 52.3
❶SATURN serves as a strong complement to math and programming. On DeepScaleR-1.5B-Preview—despite being further RL trained with 40k math and programming examples—SATURN still brings additional improvements, raising the average score from 49.9 to 52.3. Notably, it improves AIME by +5.0 and GPQA-D by +6.0. ❷ SATURN outperforms prior constructing RL task approaches across multiple models. On Qwen2.5-7B-Instruct-1M, SATURN uses only 1k training examples but improves the average score from 32.5 to 36.2, achieving a relative improvement of +8.8% over Logic-RL trained with 5k examples. These results indicate that SATURN not only complements math and programming tasks, but also provides greater improvements compared to other constructing RL task approaches.
3.6 RQ4: SATURN Enhances Self-verification in LLMs’ Reasoning Trajectories
RQ4 investigates whether SATURN affects LLMs’ reasoning behavior. On Qwen2.5-7B-Instruct-1M, we observe a gradual increase in response length during training, as illustrated in Figure 4, replicating the lengthening phenomenon reported in the R1 and Logic-RL [11, 47].
Figure 4: Response length trend during SATURN training on Qwen2.5-7B-Instruct-1M.
To examine whether such reasoning patterns generalize, we present case studies across SAT and math domains. Figure 12 shows that solving SAT variables requires rechecking all clauses, naturally encouraging self-verification. In Figure 5, SATURN-7B verifies intermediate conclusions within a small scenario and successfully chooses the correct solution path. In contrast, the baseline model reaches a wrong answer and skips verification, even when inconsistencies are detected.
Recent studies [15, 20] identify core behaviors shared by expert human reasoners and LLMs, such as verification and backtracking. These behaviors are domain-agnostic and provide fundamental reasoning patterns applicable to a wide range of tasks. In line with these findings, SATURN reinforces similar behaviors during SAT solving, leading to more structured reasoning trajectories. More reasoning trajectories are provided in Appendix J to illustrate how SATURN works. These results suggest that the self-verification patterns learned from SAT transfer well to math and programming tasks, improving reasoning robustness and reliability.
Figure 5: Comparison of reasoning trajectories on Math 500 Question No.402 between two models.
4 Discussion
4.1 Limitations of Reasoning Capability Learned from SATURN
During curriculum learning, we observed that as the number of training iterations increases, the improvements in math and programming tasks tend to plateau, which is consistent with the findings in Logic-RL. Detailed evaluation results are provided in Table 5. This plateau may stem from several factors: ❶ Knowledge limitations. SATURN improves formal logical reasoning but does not provide domain-specific knowledge supervision. This limits its effectiveness in tasks requiring mathematical or algorithmic knowledge. ❷ Context window bottlenecks. SAT problems are NP-complete tasks, and the required reasoning length grows exponentially with increasing problem difficulty. This leads to bottlenecks in the model’s capability to handle increasingly complex tasks. ❸ Limited plasticity and forgetting. Model plasticity and catastrophic forgetting are known limitations that hinder further improvements with additional training stages [1, 13].
Table 5: Performance of multi-Stage SATURN iterations on math and programming benchmarks
Model AIME 24/25 AMC 22/23 Math500 GPQA-D LiveCodeBench Avg.
DeepSeek-R1-Distill-Qwen-1.5B 21.6 65.1 83.6 30.3 16.4 43.4
SATURN-1.5B-Iteration-1 26.7 68.6 85.0 33.3 16.9 46.1
SATURN-1.5B-Iteration-2 28.3 73.5 84.6 37.4 17.4 48.2
SATURN-1.5B-Iteration-3 28.3 66.3 85.8 36.9 16.7 46.9
DeepSeek-R1-Distill-7B 50.0 80.7 93.2 49.0 35.4 61.7
SATURN-7B-Iteration-1 48.3 83.1 94.6 50.5 36.6 62.6
SATURN-7B-Iteration-2 48.3 85.5 95.0 50.5 37.7 63.4
SATURN-7B-Iteration-3 46.7 87.9 93.2 58.1 38.1 64.8
4.2 Potential of SATURN on Stronger Models
To explore the potential of SATURN to stronger models, we evaluate frontier LLMs on SAT tasks using the extended SAT instances. Results are shown in Appendix I. Although these LLMs exhibit stronger performance, they still make common errors such as hallucinating clauses, confidently committing to incorrect decisions, or failing to apply basic logical rules. Even more advanced LLMs still struggle to solve complex SAT problems. We believe SATURN remains a promising approach for enhancing reasoning in stronger LLMs. With sufficient computation, SATURN can offer a scalable, verifiable, and controllable path to further improve reasoning capabilities.
5 Related Work
5.1 Constructing Reasoning Tasks for Reinforcement Learning
Several works have explored constructing reasoning tasks to improve the reasoning capability of LLMs. Logic-RL [47] and LMRL Gym [2] train LLMs on natural language logic puzzles but lack scalability due to their limited puzzle set. ScaleQuest [12], Entity-Deducing Game [52], and K&K [46] propose automatic generation of constructing questions, but rely on LLM sampling or handcrafted templates, making large-scale generation costly and hard to verify. CodeDPO [51] and PuzzBench [41] employ LLM-based verification, which may fail silently and cannot ensure correctness. Wolf Game [45, 48] focus on multi-step logic reasoning but offer no control over task difficulty, limiting their support for curriculum learning. Overall, these tasks fall short in scalability, verifiability, or controllable difficulty. See Appendix K for detailed comparisons.
5.2 SAT-Based Evaluation of LLM Reasoning Capability
Recent studies have evaluated the reasoning capability of LLMs on SAT problems. Most of these works focus on analyzing model behavior around the SAT phase transition [18, 34, 30], where problem hardness peaks. However, the phase transition theory is originally designed for heuristic SAT solvers and does not align well with the reflective and verification-based reasoning processes of humans or LLMs. These studies also lack a fine-grained scalable difficulty framework and typically divide difficulty based on the phase transition threshold. They are further limited to supervised fine-tuning and do not consider large reasoning models with long-CoT reasoning capability trained via RL. Our work addresses these limitations by building a progressive evaluation and curriculum learning pipeline, enabling precise difficulty control and the generalization of LLMs.
6 Conclusion and Future Work
We present SATURN, a SAT-based RL framework for unleashing and evaluating the reasoning capability of LLMs. By leveraging SAT’s scalability, verifiability, and controllable difficulty, SATURN addresses key limitations of existing RL tasks. It constructs a multi-stage curriculum to gradually enhance reasoning, and introduces the SATURN-2.6k benchmark for controlled evaluation. Applied to DeepSeek-R1-Distill-Qwen, SATURN produces SATURN-1.5B and SATURN-7B, which show strong gains on unseen SAT tasks and generalize well to math and programming benchmarks.
In future work, we plan to: (1) apply SATURN to larger-scale LLMs, (2) break the existing paradigm’s reliance on human-annotated data and explore new paths toward building LLMs with continuous self-evolution capabilities.
Acknowledgements
This research is supported by the National Key R&D Program under Grant No. 2023YFB4503801, the National Natural Science Foundation of China under Grant No. 62192733, 62192730, 62192731, the Major Program (JD) of Hubei Province (No.2023BAA024), and the Beijing Natural Science Foundation under Grant No. 4264107.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the main contributions, including the curriculum learning framework SATURN and the SATURN-2.6k dataset. All supported by experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of SATURN in Section 4.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Any informal proof provided in the core of the paper is complemented by formal proofs in the appendix, including the rigorous derivation of SAT estimation.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We release all code, models, datasets, and experimental results to ensure full reproducibility.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We release all code, models, datasets, and experimental results, along with detailed information on the execution environment to ensure full reproducibility.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide all important parameters needed to understand the results in the appendix, and the code scripts include complete training and test details.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Following current evaluation practices in recent large model benchmarks, we report the unbiased estimator of the pass@k metric. This expectation-based metric is designed to reflect statistical significance across multiple samples and inherently captures sampling variability. Although we do not report explicit error bars, pass@k inherently reflects statistical reliability. Further details are provided in Appendix E.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide details on the GPU resources used for experiments as well as information on all data storage.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed and ensured that all research conducted in this paper fully conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper focuses on improving LLMs’ reasoning capability through SATURN curriculum learning framework, without direct deployment or application, and thus does not raise societal impacts concerns.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no foreseeable risks of misuse, as it does not involve the release of high-risk models or datasets.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We properly credit all used assets in the paper and appendix, including code, datasets, and models. Licenses and terms of use are explicitly stated and respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release new code, datasets, and models. All new assets are thoroughly documented, including usage instructions and data format specifications, which are provided alongside the release.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve any form of crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing or research with human subjects, so IRB approval is not applicable.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: This paper does not involve the use of LLMs as an important, original, or non-standard component of the core methodology.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix
  • Appendix A: Pseudocode of SATURN Algorithm
  • Appendix B: Procedure of SAT Constructor
  • Appendix C: Derivation and Validity of the SAT Difficulty Estimation
  • Appendix D: Training Schedule for SATURN-1.5B and SATURN-7B
  • Appendix E: Evaluation Hyperparameters
  • Appendix F: Prompt Templates
  • Appendix G: Detailed Performances of SATURN Models on SATURN-2.6k
  • Appendix H: Ablation Details for SATURN
  • Appendix I: Behavior of Stronger LLMs on SATURN Tasks
  • Appendix J: Examples of LLMs Reasoning Trajectories
  • Appendix K: Comparisons of Existing Reasoning Tasks
  • Appendix L: SATURN-7B Word Cloud on GPQA Diamond
Appendix A Pseudocode of SATURN Algorithm and Hyperparameters
Algorithm 1 presents the complete algorithmic workflow of SATURN.
[…]
def Increase_difficulty(n, k, l, step=1):
Increments by D_step to increase SAT difficulty.
return n, k, l + D_step
def SATURN_learning_loop(n, k, l, LLM):
for t in range(2): # Max total curriculum iterations
# Step 1: Curriculum Estimation Loop
# Generate validation set
Val_set = SAT_Construction(n, k, l, Val_size)
pass_at_1 = evaluate(Val_set, LLM) # Evaluate pass@1
if pass_at_1 >= epsilon:
n, k, l = Increase_difficulty(n, k, l)
else:
# Step 2: LLM Training Loop (only if pass_at_1 < epsilon)
for i in range(10): # Max GRPO steps per level
# Generate training set
D_train = SAT_Construction(n, k, l, Train_size)
LLM = GRPO(LLM, D_train)
# Re-generate validation set
Val_set = SAT_Construction(n, k, l, Val_size)
pass_at_1 = evaluate(Val_set, LLM) # Evaluate pass@1
if pass_at_1 >= epsilon:
break
return LLM
Table 6 shows all SATURN hyperparameters.
Table 6: SATURN Hyperparameters
Parameter SATURN-1.5B SATURN-7B
Initial $(n,k,l)$ (3, 5, 5) (3, 5, 13)
pass@k threshold $\epsilon$ 0.5 0.75
Training set size per step (Train_size) 250 250
Validation set size per step (Val_size) 40 40
Difficulty increment (D_step) 1 2
Curriculum iterations 2 2
Max GRPO steps per level 10 10
Appendix B Construction Procedure of SAT Instances
This appendix describes the implementation details of the SAT_Construction algorithm introduced in Section 2.1. The goal is to generate $m$ satisfiable $(n,k,l)$-SAT instances in conjunctive normal form (CNF), each constructed to be consistent with a known Boolean solution. The algorithm ensures diversity and uniformity across sampled instances.
The construction procedure is outlined in Algorithm 2 and consists of the following steps:
  1. 1.
    A Boolean solution is randomly generated for the $k$ variables.
  2. 2.
    The constructor randomly selects $n$ variables from the $k$ variables to form a clause.
  3. 3.
    Under the constraint of satisfying the solution, the $n$ variables are randomly negated, resulting in $2^{n}-1$ satisfiable clauses per variable set. From these steps, we uniformly obtain a large number of single-clause samples from the total of $2^{k}\cdot\binom{k}{n}\cdot(2^{n}-1)$ valid $(\texttt{solution},\texttt{clause})$ pairs, with an upper bound set to 100,000100{,}000.
  4. 4.
    These $(\texttt{solution},\texttt{clause})$ pairs are grouped into clusters based on the same solution.
  5. 5.
    Within each cluster, we randomly select $l$ clauses and shuffle their order to construct a full SAT instance that satisfies the corresponding solution.
  6. 6.
    Finally, we uniformly sample across clusters and remove duplicates to obtain a total of $m$ diverse SAT instances.
Algorithm 2 SAT_Construction(n,k,l,m)(n,k,l,m)
def SAT_Construction(n, k, l, m):
V = [x_1, x_2, ..., x_k] # k Boolean variables
P = set() # pool of (solution, clause) pairs
# Step 1-3: Generate (solution, clause) pairs
while len(P) < 100000: # max number of single-clause samples
vars = sample_variables(V, n) # select n vars from V
# e.g., [x_2, x_4, ...]
solution = random_assign(vars) # assign 0/1 to vars
# e.g., {x_2:1, x_4:0, ...}
clause = generate(vars, solution) # create clause
# e.g., {x_2 or !x_4 ...}
P.add((solution, clause)) # store pair
# Step 4: Group by solution
clusters = group_by_solution(P) # solution -> list of clauses
# Step 5-6: Construct m SAT instances
instances = set()
while len(instances) < m:
solution, clauses = sample_cluster(clusters) # select a group
if len(clauses) < l:
continue
selected_clauses = sample_clauses(clauses, l) # pick l clauses
shuffle(selected_clauses)
instances.add((solution, selected_clauses))
return instances
Appendix C Derivation and Validity of Difficulty Estimation for SAT Tasks on LLMs
As stated in Section 2.2, we propose a composite difficulty function Eq. (2) to estimate the reasoning difficulty of SAT problems for LLMs. This difficulty score combines a sparsity-based estimate of solution density with a structural complexity adjustment.
Step 1: Sparsity-Based Estimate ($D_{1}$)
We first estimate task difficulty by measuring the ratio between symbolic search cost and expected solution space size.
The symbolic search cost is approximately proportional to the number of variable-symbol combinations across decoding steps. For a problem with $k$ variables and $l$ decoding steps, we estimate:
$\mathbf{Search\ Cost}\propto k\cdot l$ (4)
For Boolean constraint problems like SAT, the number of satisfying assignments is sparse. Assuming a random clauses, the expected number of valid solutions is:
$\mathbf{Expected\ Solutions}\approx 2^{n}-1\approx 2^{n}$ (5)
where we approximate $2^{n}-1$ by $2^{n}$ for analytical simplicity.
Taking the ratio and applying a logarithmic transform yields:
$D_{1}(n,k,l)=\log_{2}(k\cdot l)-\log_{2}(2^{n})=\log_{2}(k)+\log_{2}(l)-n$ (6)
Step 2: Structural Complexity Adjustment ($D_{2}$)
In addition to solution sparsity, symbolic reasoning difficulty also depends on the structural properties of the input formula. We construct a structure-aware term $D_{2}(n,k,l)$ based on two contributing factors.
First, consider the symbolic reuse density: when $k$ variables are reused across $n$ clauses, each variable is, on average, involved in $n/k$ constraints. This increases the interdependency between clauses. Since higher reuse leads to greater symbolic entanglement, making factorization more challenging for the model, we define the inverse ratio $\frac{k}{n}$ as a proxy for the structural cost:
$\mathbf{Variable\ Interaction\ Cost}=\frac{k}{n}$ (7)
Second, the clause length $l$ determines the number of symbols that each clause contains. Longer clauses introduce more intra-clause dependencies, increasing local reasoning complexity. We approximate this with:
$\mathbf{Clause\ Width\ Cost}=\log_{2}(l)$ (8)
These two components affect reasoning difficulty independently—one globally through variable sharing, and the other locally through clause complexity. We, therefore, combine them additively into:
$D_{2}(n,k,l)=\frac{k}{n}+\log_{2}(l)$ (9)
This additive form is justified as both terms grow monotonically with symbolic complexity and are approximately aligned in scale, enabling stable and interpretable difficulty estimation.
Step 3: Final Composite Metric
Combining both components yields:
$D(n,k,l)$ $=D_{1}(n,k,l)+D_{2}(n,k,l)$ (10)
$=\log_{2}(k)+\log_{2}(l)-n+\frac{k}{n}+\log_{2}(l)$ (11)
$=\log_{2}(k)+2\log_{2}(l)-n+\frac{k}{n}$ (12)
This final composite metric provides a stable and interpretable approximation of symbolic difficulty for LLMs, taking into account both sparsity and structure.
Step 4: Ablation Study about estimation metric
To further validate the necessity and effectiveness of Eq. (12) composite metric, we conduct an ablation study comparing alternative formulations. Table 7 reports the $R^{2}$ correlation of each metric with LLM performance (pass@3) across multiple model scales.
Table 7: Metric comparison across models.
Metric Formula R1-Qwen-1.5B Saturn-1.5B R1-Qwen-7B Saturn-7B Avg. Std. Dev.
$k-l\cdot\log_{2}(1-\frac{1}{2^{n}})$ 0.507 0.537 0.132 0.000 0.294 0.269
$\alpha\cdot\frac{k}{n}+\beta\cdot\log_{2}(l),\ \alpha=2,\ \beta=1$ 0.428 0.478 0.568 0.508 0.496 0.059
$\alpha\cdot\frac{k}{n}+\beta\cdot\log_{2}(l),\ \alpha=1,\ \beta=1$ 0.240 0.279 0.719 0.826 0.516 0.300
$\log_{2}(k\cdot l)-\log_{2}(2^{n}-1)$ 0.875 0.893 0.451 0.157 0.594 0.356
$\log_{2}(k)+2\cdot\log_{2}(l)-n+\frac{k}{n}$ 0.707 0.746 0.724 0.501 0.670 0.113
Simpler metrics that consider only sparsity or only structural complexity perform worse overall, confirming that both components are essential for accurately capturing task difficulty. Our metric $\log_{2}(k)+2\cdot\log_{2}(l)-n+\frac{k}{n}$ combines both solution sparsity and structural complexity, which achieves the best overall correlation. The ratio between the solution space and the LLM’s search space is the most crucial aspect. On SATURN‑7B, the $R^{2}$ value is about 0.5 because the LLM already achieves over 90% accuracy on easy problems, which limits the observable linear correlation in that range. The corresponding figures are shown in Figures 10 and 11.
Appendix D Training Schedule for SATURN-1.5B and SATURN-7B
This appendix provides additional details on the SATURN-1.5B and SATURN-7B training. We conduct all experiments on 8 NVIDIA A100 40GB GPUs. We use the OpenRLHF framework [^1] https://github.com/OpenRLHF/OpenRLHF for GRPO training. The framework is designed to make RL training simple and user-friendly, which works well in our experiments. The hyperparameters used in training are summarized in Table 8. All other parameters not listed above consistently follow the default settings of OpenRLHF.
Table 8: OpenRLHF Training Hyperparameters
Parameter SATURN-1.5B SATURN-7B
Actor learning rate $5\times 10^{-7}$ $5\times 10^{-7}$
Initial KL coefficient $\times 10^{-3} 1×1$$\times 10^{-3}$
Batch size (train) 2 2
Batch size (rollout) 2 2
Samples per prompt 8 8
Prompt length (max) 1024 1024
Generation length (max) 10000 8192
Temperature 0.8 1.0
Zero redundancy stage 3 3
Use bf16 Yes Yes
KL estimator k3 k3
Advantage estimator GroupNorm GroupNorm
Use reward normalization Yes Yes
Figure 6 and Figure 7 illustrate the training curves of SATURN-1.5B and SATURN-7B. Both models show a clear upward trend in the average reward during training, accompanied by early fluctuations that gradually stabilize. The maximum reward curves quickly reach high values and maintain them throughout most of the training process. These results indicate that both SATURN-1.5B and SATURN-7B successfully learn to generate high-reward outputs, demonstrating effective SAT reward-guided optimization.
Figure 6: Training curves of various metrics for SATURN-1.5B.
Figure 7: Training curves of various metrics for SATURN-7B.
Appendix E Evaluation Hyperparameters
This appendix provides additional details on the hyperparameters used in the evaluation. We use the Hugging Face lighteval library [^2] https://github.com/huggingface/lighteval for math and programming evaluations. It offers efficient benchmarks, helping us assess LLMs’ performance across various tasks while ensuring both computational efficiency and high-quality results. For the evaluation of the DeepSeek-R1 series distillation models, we use lighteval with the Hugging Face-Open-R1 framework [^3] https://github.com/huggingface/open-r1. This framework effectively reproduces the evaluation results of the R1 series models.
During the evaluation process, we follow the parameter settings from Hugging Face-Open-R1, as shown in Table 9. Additionally, for LiveCodeBench, we also select the default v4_v5 version of this framework. Due to the long inference budget required by LiveCodeBench, we set the maximum response length to 16K and generate four samples per instance to estimate pass@1. All other parameters not listed and mentioned above consistently follow the default settings of Hugging Face-Open-R1.
For larger closed-source models, we report the benchmark results from a public benchmark website [^4] https://artificialanalysis.ai/models.
Table 9: Evaluation Hyperparameters for Hugging Face-Open-R1
Hyperparameter Setting
Data type bfloat16
Maximum model length 32,768
Maximum new tokens 32,768
Temperature 0.6
Top-$p$ (nucleus sampling) 0.95
Appendix F Prompt Templates
This appendix provides the prompt templates used for evaluation, ensuring consistency and reproducibility across tasks. Figure 8 presents the format for SAT problem training and evaluation, while Figure 9 shows the template used for math, programming, and GPQA Diamond tasks.
Figure 8: Prompt format used for SAT problem training and evaluation.
Figure 9: Prompt format used for math, programming, and GPQA Diamond problems evaluation.
Appendix G Detailed Performances of SATURN models on SATURN-2.6k
This appendix provides additional details of SATURN-1.5B and 7B on SATURN-2.6k spanning 10 harder SAT difficulty levels. Experimental results are shown in Table 1014, and Figures 1011. We summarize two key observations:
❶The pass@3 accuracy correlates strongly with the estimated SAT difficulty $D(n,k,l)$ across models. Specifically, the linear regression $R^{2}$ scores are 0.746 for SATURN-1.5B and 0.707 for DeepSeek-R1-Distill-Qwen-1.5B (Figure 10), and 0.5011 for SATURN-7B and 0.724 for DeepSeek-R1-Distill-Qwen-7B (Figure 11). These results indicate that our difficulty function effectively captures problem hardness, supporting the design of a curriculum learning schedule based on it. They also demonstrate that SAT is a reliable benchmark for evaluating reasoning capability.
❷Although our models are trained only on relatively easier SAT problems, they show consistent improvements on harder levels. As shown in Table 1014, both SATURN-1.5B and SATURN-7B generalize well to more challenging problems, highlighting the effectiveness of our curriculum-driven training strategy.
Table 10: Full pass@1 results on SATURN-2.6k
Model (3,7,40) (3,5,25) (3,5,20) (3,6,20) (3,7,20) (4,7,40) (4,8,40) (4,7,20) (6,7,40) (5,8,40) Avg.
R1-Distill-1.5B 0.3 1.3 3.3 3.3 3.3 1.1 1.6 7.2 10.6 4.4 3.6
SATURN-1.5B 0.3 6.2 7.8 7.3 8.9 3.3 4.2 22.5 29.0 12.9 10.2
R1-Distill-7B 2.3 12.3 19.2 23.3 21.5 3.8 4.5 30.1 24.4 12.3 15.4
SATURN-7B 8.6 44.7 66.4 64.5 64.6 9.2 10.3 57.3 36.4 19.2 38.1
Table 11: Full pass@3 results on SATURN-2.6k
Model (3,7,40) (3,5,25) (3,5,20) (3,6,20) (3,7,20) (4,7,40) (4,8,40) (4,7,20) (6,7,40) (5,8,40) Avg.
R1-Distill-1.5B 1.0 3.8 9.4 9.1 9.3 3.3 4.7 19.3 28.9 12.5 10.1
SATURN-1.5B 0.8 15.8 19.5 18.5 21.6 9.5 11.3 50.3 62.5 31.7 24.2
R1-Distill-7B 6.5 31.2 45.6 52.8 49.9 10.4 11.8 65.6 56.2 31.3 36.1
SATURN-7B 22.9 78.2 94.1 93.5 93.3 24.3 26.5 91.0 72.4 45.7 64.2
Table 12: Full pass@5 results on SATURN-2.6k
Model (3,7,40) (3,5,25) (3,5,20) (3,6,20) (3,7,20) (4,7,40) (4,8,40) (4,7,20) (6,7,40) (5,8,40) Avg.
R1-Distill-1.5B 1.7 6.1 14.9 14.2 14.4 5.4 7.6 29.2 43.7 19.6 15.7
SATURN-1.5B 1.3 23.5 27.9 26.7 30.6 14.9 17.2 67.0 79.0 44.1 33.2
R1-Distill-7B 10.5 44.7 62.3 69.2 66.9 16.0 17.5 82.9 74.0 45.4 48.9
SATURN-7B 34.4 89.5 98.5 98.6 98.2 35.9 38.6 97.5 86.4 62.4 74.0
Table 13: Full pass@7 results on SATURN-2.6k
Model (3,7,40) (3,5,25) (3,5,20) (3,6,20) (3,7,20) (4,7,40) (4,8,40) (4,7,20) (6,7,40) (5,8,40) Avg.
R1-Distill-1.5B 2.3 8.1 19.9 18.9 18.9 7.6 10.1 37.4 55.7 25.8 20.5
SATURN-1.5B 1.8 30.1 34.2 32.8 37.5 19.8 22.1 78.2 87.8 52.5 39.7
R1-Distill-7B 14.3 54.6 73.6 78.7 77.8 20.9 22.2 91.6 84.3 56.3 57.4
SATURN-7B 43.7 94.2 99.6 99.7 99.5 45.0 47.6 99.2 92.2 73.5 79.4
Table 14: Full pass@10 results on SATURN-2.6k
Model (3,7,40) (3,5,25) (3,5,20) (3,6,20) (3,7,20) (4,7,40) (4,8,40) (4,7,20) (6,7,40) (5,8,40) Avg.
R1-Distill-1.5B 3.3 10.6 26.8 25.2 24.6 10.8 14.5 47.9 69.3 33.7 26.7
SATURN-1.5B 2.5 38.7 41.3 39.5 45.7 26.2 28.0 89.6 101.0 75.9 46.7
R1-Distill-7B 19.6 65.0 84.7 86.6 87.8 27.2 27.9 97.4 92.8 68.6 65.8
SATURN-7B 54.9 97.1 99.9 99.9 100.0 55.0 57.3 99.8 95.3 84.7 84.4
Figure 10: Scatter plots of pass@3 versus estimated difficulty $D(n,k,l)$ for DeepSeek-R1-Distill-Qwen-1.5B and SATURN-1.5B, with linear regression fits. The linear regression for two models achieve $R^{2}$ values of 0.707 and 0.746 respectively.
Figure 11: Scatter plots of pass@3 versus estimated difficulty $D(n,k,l)$ for DeepSeek-R1-Distill-Qwen-7B and SATURN-7B, with linear regression fits. The linear regression for two models achieve $R^{2}$ values of 0.724 and 0.5011 respectively.
Appendix H Ablation Studies for SATURN
This appendix presents the ablation studies for SATURN, as shown in Table 15 and Table 16. Each training setting is denoted as $(n,k,l)\times\texttt{Train\_size}$, where $(n,k,l)$ controls SAT construction and Train_size is the number of training examples. Here, Train_size can also be written as Train_size = $D\times\texttt{num}$, where $D$ is the number of difficulty levels and num is the number of samples per level in SATURN-2.6k (Figure 11 and 10). These experiments validate the effectiveness of curriculum learning and the design of various SAT training configurations.
Table 15: Ablation comparison on math and programming benchmarks
Training Setting AIME 24/25 AMC 22/23 Math500 GPQA-D LiveCodeBench Avg.
$(3,5,10)\times 500$ 45.0 84.3 93.2 53.5 35.1 62.2
$(3,5,13)\times 500$ 50.0 83.1 94.6 50.0 36.1 62.8
$(3,5,15)\times 500$ 35.0 68.7 86.6 46.0 31.3 53.5
$(3,5,13)\times 1000$ 51.7 81.9 94.0 47.5 37.2 62.5
$((3,5,13)+(3,5,15))\times 500$ (one epoch) 43.3 86.7 93.0 49.5 35.8 61.7
$(3,5,13)\times 500+(3,5,15)\times 500$ 48.3 85.5 95.0 50.5 37.7 63.4
Table 16: Ablation Study with Different Sampling Strategies and Training Budgets
Training Setting AIME 24/25 AMC 22/23 Math500 GPQA-D LiveCodeBench Avg.
$(n,k,l)\times 100\times 10$ + shuffle 38.3 66.3 90.6 44.9 35.2 55.1
$(n,k,l)\times 100\times 10$ 46.7 85.5 93.2 50.5 36.9 62.3
$(n,k,l)\times 200\times 5$ + shuffle 48.3 81.9 93.0 52.0 35.8 62.2
$(n,k,l)\times 200\times 5$ 46.7 88.0 93.2 48.0 35.8 62.6
$(3,5,13)\times 500+(3,5,15)\times 500$ 48.3 85.5 95.0 50.5 37.7 63.4
In Table 15, we evaluate the impact of SAT difficulty, training budgets, and curriculum structure. We draw two key conclusions:
SATs that are too easy or too hard hinder model learning. Training solely on easy $(3,5,10)\times 500$ or hard $(3,5,15)\times 500$ instances results in lower average scores (62.2 and 53.5, respectively). In contrast, moderate-difficulty SATs $(3,5,13)\times 500$ yield a higher score of 62.8, showing that balanced difficulty is essential for effective reasoning development.
Multi-stage curriculum learning outperforms flat or mixed training. Curriculum learning with progressively increasing SAT difficulty $(3,5,13)\times 500+(3,5,15)\times 500$ achieves the highest average score of 63.4. In contrast, one-epoch mixed training $((3,5,13)+(3,5,15))\times 500$ only reaches 61.7, despite using the same total number of examples. Furthermore, simply scaling up a single-stage setting $(3,5,13)\times 1000$ yields 62.5, which is also inferior to the curriculum. These results indicate that progressive difficulty scheduling is more effective than either flat or mixed training with the same or larger data budget.
Table 16 further investigates the impact of training thresholds and step sizes under a fixed total training budget.
Gradual difficulty progression outperforms random shuffling of difficulty levels. Both $(n,k,l)\times 100\times 10$ and $(n,k,l)\times 200\times 5$ perform better when difficulty levels follow a gradual progression (62.3 and 62.6), compared to random shuffling of difficulty levels (55.1 and 62.2). This demonstrates that a curriculum learning approach with progressive difficulty scheduling is more effective.
Excessively fine-grained difficulty levels hinder performance. Training with overly fine-grained difficulty levels, such as $(n,k,l)\times 100\times 10$, results in lower performance (55.1) compared to coarser steps like $(n,k,l)\times 200\times 5$ (62.6). Both of these configurations perform worse than the two-stage curriculum $(3,5,13)\times 500+(3,5,15)\times 500$, which achieves the highest performance with an average score of 63.4. This indicates that excessively fine-grained difficulty levels prevent the model from effectively mastering each level before moving on to the next, hindering overall learning.
Appendix I Behavior of Stronger LLMs on Extended SATURN Tasks
This appendix demonstrates the performance of stronger LLMs on more challenging SAT tasks. The experimental results are shown in Table 17. Even the strongest LLMs available today cannot solve complex SAT tasks as effectively as humans using simple reflection and verification search. Due to the long CoT involved, the full LLMs’ outputs are provided in the supplementary material.
Table 17: One-shot performance of stronger LLMs on extended SATURN tasks. Models with ✔ successfully solve the corresponding difficulty of SAT tasks. Kimi-1.5 solves the task but with significantly longer reasoning chains.
SAT Task (n, k, length) GPT-4o O1-mini DeepSeek-V3 R1 Kimi-1.5
(3, 5, 30) ✔*
(4, 7, 80)
To provide a baseline comparison, we also tested a CDCL SAT solver  [^5] https://en.wikipedia.org/wiki/Conflict-driven_clause_learning on the SATURN-2.6k test set. The results are as follows:
Table 18: CDCL SAT solver performance on SATURN-2.6k test set.
Metric Value
Total instances 1000
Satisfiable 1000
Unsatisfiable 0
Valid SAT Models 1000
Model Accuracy 100.00%
Total time taken (s) 0.14
Table 18 shows that the CDCL SAT solver significantly outperforms current LLMs like DeepSeek-R1, both in terms of runtime and accuracy. SATURN applied to today’s strongest LLMs still has great potential.
Appendix J Examples of Different LLMs’ Reasoning Trajectories
This appendix presents examples of different LLMs’ reasoning trajectories, as shown in Figures 5 and 12. The full LLMs’ outputs are provided in the supplementary material.
Figure 12: Case study of SATURN-7B’s reasoning trajectory on a SAT instance.
Figure 13: Comparison of reasoning trajectories on Math 500 Question No.41 between two models.
In the case of Math500-41 (Table 13), SATURN improves its ability to avoid unnecessary calculations and dead-end reasoning paths. By leveraging the self-verification patterns learned from SAT tasks, such as "I made a mistake earlier," SATURN-7B can better navigate through the problem, discarding invalid paths earlier in the process.
In the case of Math500-402 (Table 5), SATURN-7B demonstrates a stronger ability to self-verify intermediate conclusions in a smaller scenario (with 4 individuals). While DeepSeek-R1-Distill-Qwen-7B also tries to identify a smaller scenario, it fails to recheck the result when an inconsistency is found, instead stating, "perhaps I’m overcomplicating this." In contrast, SATURN-7B can identify the error and re-verify the results within this small scenario, ultimately selecting the correct solution from two possible candidates.
In conclusion, SATURN-7B exhibits enhanced self-verification capabilities. LLMs sometimes confidently claim that a wrong answer is correct. Solving SAT tasks inherently involves frequent and fine-grained clause verification, which trains LLMs to perform precise checking during reasoning. The self-verification patterns learned from SAT tasks help LLMs solve math problems more effectively by selecting correct solutions from multiple options. These results suggest that the self-verification mechanisms developed during SAT learning (Table 12) generalize well to math and programming tasks, improving the LLMs’ reasoning robustness and reliability.
Appendix K Detailed Comparison of Constructed Reasoning Tasks
Figure 14: Comparison of Knights and Knaves (K&K) logic puzzle and SAT problem.
This appendix provides a detailed comparison between K&K logic puzzle and SAT problem. Illustrative examples of each type are shown in Figure 14. While both are designed to test reasoning capability, they differ in construction cost, output format complexity, and difficulty control.
Construction Cost.
K&K puzzles require translating symbolic logic into natural language. This involves paraphrasing logical constraints into grammatically and semantically valid sentences, which increases construction cost. In contrast, SAT problems are purely symbolic and follow a standard format. As shown in [18], symbolic SAT and its natural language version yield similar reasoning performance for LLMs, suggesting the symbolic form is sufficient.
Output Format Complexity.
[…]
Difficulty Control.
K&K puzzles use the number of characters to control difficulty, which is coarse-grained. SAT problems allow fine-grained control via clause structure and variable interactions. We further define an estimation of SAT task difficulty for LLMs as $D(n,k,l)=\log_{2}(k)+2\log_{2}(l)-n+\frac{k}{n}$. Adding a clause to a SAT formula never decreases its difficulty, for both humans and LLMs. This makes SAT more suitable for curriculum learning.
In summary, while K&K puzzles provide linguistic diversity, SAT problems are more efficient in construction, output consistency, and difficulty regulation, making them preferable for training LLMs at scale.
Appendix L Word Cloud of SATURN-7B’s Outputs on GPQA Diamond
Figure 15: Word cloud of SATURN-7B’s generated answers on GPQA Diamond. Frequently used tokens are shown in larger fonts.
Figure 15 shows the word cloud of SATURN-7B’s generated answers on GPQA Diamond, highlighting its frequent use of self-verification patterns in reasoning.